OpenStudy (jkbo):

What is the solution to the system of equations? -x-y-2z=-2 -x-y+z=5 3x+2y-z=-1 a.(4, –4, –5) b.(4, –4, 5) c.(4, 4, 5) d.(–4, –4, 5)

OpenStudy (anonymous):

\[\left\{x=\frac{20}{3},y=-\frac{28}{3},z=\frac{7}{3}\right\} \]

OpenStudy (jkbo):

Thank you but Im not sure what to do from here

OpenStudy (jkbo):

@robtobey

OpenStudy (anonymous):

a.(4, –4, –5) b.(4, –4, 5) c.(4, 4, 5) d.(–4, –4, 5) All are incorrect answers. From Mathematica:\[\{-x-y-2 z= -2\}\text{/.}\{ x\to 4,y\to -4,z\to 5\} \to \{\text{False}\}\]

OpenStudy (anonymous):

The fractions I posted are the correct answers.

OpenStudy (jkbo):

ok thank you but Those Are the only answers I have.

OpenStudy (anonymous):

all are incorrect

OpenStudy (jkbo):

sorry for the first one its x - y - 2z = -2

OpenStudy (jkbo):

@rkreddy1986

OpenStudy (anonymous):

With the change, x - y - 2z = -2, \[\{x=4,y=-4,z=5\} \] are correct.

OpenStudy (jkbo):

Thank you